# AgentExpt: Automating AI Experiment Design with LLM-based Resource Retrieval Agent

### v1

@@ 4.2. Automated Data Collection Pipeline
We systematically acquired research papers and corresponding metadata published between 2015 and 2025 from prominent AI academic repositories, including OpenReview (Ope, [n. d.]), arXiv (arX, [n. d.]), and official conference websites of ten leading AI venues. The collected metadata includes comprehensive information such as paper titles, authors, abstracts, publication years, conference affiliations, and direct links to publications. After rigorous deduplication and cleaning, the corpus exceeding one hundred thousand papers was refined into a high-quality dataset suitable for subsequent analyses. Utilizing the MinerU tool (Wang et al., 2024), we performed structured parsing on each paper, accurately extracting chapter structures, paragraph boundaries, table arrangements, and citation anchors, thereby ensuring reliable downstream entity extraction.
@@
To identify baseline methods and datasets used in experiments, we analyzed citations within experiment-centric sections (e.g., Experiments, Results, Evaluation) as primary anchors. Baseline methods were extracted by analyzing citation contexts and adjacent text, with additional metadata, such as introductory papers and repository links, collected from resources like Papers with Code (PWC) (PWC, [n. d.]) and OpenAlex (Priem et al., 2022). For dataset identification, we used the DataStet (Lopez, 2025) model to automatically extract dataset entities from the experimental sections, validating them through metadata from platforms like Hugging Face (Hug, [n. d.]) and Kaggle (Kag, [n. d.]). This process included entity normalization, alias resolution, and filtering, ensuring high-quality and accurate data for subsequent validation stages.
Figure 5. Illustration of Collective Perception Augmented Retrieval:
@@
To further enhance extraction accuracy and effectively mitigate false-positive candidates, we introduced a two-stage filtering strategy. Initially, rule-based filtering mechanisms were formulated to preliminarily evaluate baseline and dataset candidates by leveraging multiple criteria, such as contextual positioning, citation frequency, and naming consistency. Subsequently, entities situated near decision threshold boundaries underwent additional verification through few-shot prompting using large language models (LLMs). Specifically, baseline entities were subjected to LLM-based context-aware evaluation, integrating external metadata, while dataset entities initially underwent metadata-driven normalization and deduplication, followed by high-confidence LLM-based verification combining textual context and metadata. This dual-stage filtering strategy substantially improved overall dataset accuracy and reliability.
4.3. Data Analysis
We conducted an extensive quality assessment and statistical analysis of the finalized scholarly dataset. This comprehensive dataset comprises 108,825 research papers published between 2015 and 2025 across ten leading artificial intelligence conferences, associated with 116,970 distinct baseline methods and 68,316 evaluation datasets. On average, each paper is linked to 5 baselines and 3 datasets, with the completeness ratio of repository links achieving 13%.
A key finding from our analysis is the remarkably high reuse rate of established resources within the community. As illustrated in Figure 4, both baselines and datasets exhibit strong recurring adoption patterns over time. For instance, the baseline reuse rate consistently remains above 70% in recent years, with several periods even exceeding the 90% threshold. This indicates that our dataset successfully captures the core set of canonical methods and benchmarks that form the foundation for experimental comparisons in AI research, rather than just a superficial collection of mentioned names.
@@ 4.3. Data Analysis
5. Method
5.1. Overview
We aim to recommend experimentally suitable baselines and datasets for a target research idea, rather than merely retrieving items with high topical similarity. Most prior approaches (Viswanathan et al., 2023b; Färber and Leisinger, 2021a, b; Beltagy et al., 2019; Qayyum et al., 2025; Goyal et al., 2024; Gu et al., 2022) represent each candidate only from a first–person perspective (its self–description in the introducing paper or repository card). Such views often diverge from how the community actually positions and uses the candidate, omitting crucial third–person signals from citation and usage contexts. In addition, baseline and dataset recommendation are typically treated independently, whereas in real workflows the two are tightly coupled: when selecting datasets we routinely consider those that co–occur with plausible baselines (and vice versa), reflecting a strong baseline–dataset synergy.
Motivated by these observations (Shen and Barabási, 2014), we propose a two–stage recommendation framework. (1) Collective Perception–Augmented Retrieval. We represent each candidate by concatenating its self–description with collective perception, and finetune an embedding model on these representations to improve recall of components. (2) Reasoning–Augmented Reranking. We extract paper-baseline–paper-dataset(paper-dataset–paper-baseline) interaction chains from the literature and train a language model to generate reasoning chains over these interactions to rerank the shortlist.
5.2. Citation Context Augmented Retrieval
This stage builds retrieval representations that jointly encode a candidate’s first–person self–description and third–person collective perception distilled from citation contexts, and then finetunes a dense retriever to maximize recall of baselines and datasets, see fig 5.
@@ 5.2. Citation Context Augmented Retrieval
Our goal is to capture a target’s collective perception. A naive choice is to treat the pooled citation contexts $C(t)$ as this signal. We avoid this because a popular target may accrue thousands of mentions, far beyond the embedding model’s context window. To address these issues, We instantiate a target’s Collective Perception(CP) as a citation context summary. For each target $t$, we use GPT4o (OpenAI, 2024) to synthesize $\mathrm{CP}(t)$ from its pooled citation contexts $C(t)$. The summary answers when and why the baseline or dataset is used, distilling recurring signals about task assumptions, evaluation protocols (metrics, splits, scales), typical configurations, and common co-usage patterns. During synthesis, we de-duplicate paraphrases, discard background or non-experimental mentions, and enforce an evidence-grounded prompt that favors majority-supported observations and avoids speculation.
@@
where $\mathcal{T}$ describes the task (e.g., “Given a research idae, retrieve relevant baseline methods or datasets that are most suitable.”). Candidates remain in their native form without an instruction:
@@
$\mathcal{L}=\underbrace{\frac{1}{B}\sum_{i=1}^{B}\mathrm{BCE}\!\big(S_{ii},\,y_{i}\big)}_{\text{diagonal (main)}}\;+\;\lambda\,\underbrace{\frac{1}{B(B-1)}\sum_{i\neq j}\mathrm{BCE}\!\big(S_{ij},\,0\big)}_{\text{in-batch contrastive}}.$
@@
5.3. Reasoning–Augmented Reranking
This stage constructs reasoning chain representations from observed P-B-P-D(P-D-P-B) interactions and finetunes a language model to rerank the shortlist with interpretable, chain-based justifications.
@@ 5.3. Reasoning–Augmented Reranking
which supports baseline recommendation via $p\!\to\!d\!\to\!p^{\prime}\!\to\!b$. Dataset recommendation is symmetric: $p\!\to\!b\!\to\!p^{\prime\prime}\!\to\!d$. For each hop we attach evidential snippets.
@@
To unlock LLM reasoning and enable reasoning reranking based on interaction chains, we use GPT4o (OpenAI, 2024) to generate reasoning Chain trajectories. For each query and each candidate baseline $b$, we enumerate chains of the form $p\!\to\!d\!\to\!p^{\prime}\!\to\!b$ and score each chain by
@@
$m_{B\!\to D}(b,d)\;=\;\bigl|\{\,p^{\prime}\;:\;p^{\prime}\text{ uses }b\text{ and }d\,\}\bigr|$
@@
Table 2. Comparison on paper–baseline and paper–dataset recommendation with conservatively estimated Recall@10/20/30 and Hitrate@5/10/15.
Paper–Baseline Paper–Dataset
Method R@10 R@20 R@30 HR@5 HR@10 HR@15 R@10 R@20 R@30 HR@5 HR@10 HR@15
6. Experiment
6.1. Experimental Setup
@@
We compare our method against five representative families: (i) the lexical matching baseline BM25 (Robertson et al., 2009); (ii) the graph/contrastive approach Textual-GCL (Qayyum et al., 2025); (iii) the scientific-domain encoder SciBERT (Beltagy et al., 2019); (iv) the attention-based model HAtten (Gu et al., 2022); and (v) the structure-aware method SymTax (Goyal et al., 2024). These cover lexical, dense, graph/contrastive, and attention/structure-aware paradigms commonly used for scientific retrieval and recommendation.
@@
[…]
6.2. Main Result
On both baseline and dataset recommendation, our method attains the best results in the main comparison. On the baseline side, our retriever improves Recall@20 from 0.4218 (SymTax) to 0.4523 (+0.0305, +7.23%), and the full two-stage model further lifts precision-oriented metrics: HitRate@5 from 0.5476 to 0.5933 (+0.0457, +8.35%) and HitRate@10 from 0.6453 to 0.6938 (+0.0485, +7.52%). Relative to BM25, Recall@20 rises from 0.2971 to 0.4523 (+52.2%). On the dataset side, our retriever increases Recall@20 over SymTax from 0.2873 to 0.3001 (+0.0128, +4.46%), while the full model reaches HitRate@5 0.4557 vs. 0.4210 (+0.0347, +8.24%) and HitRate@10 0.5549 vs. 0.5125 (+0.0424, +8.27%). Compared to BM25, Recall@20 improves from 0.1906 to 0.3001 (+57.5%). As expected, reranking the top-20 does not change Recall@20 but consistently boosts HitRate, especially for datasets (from 0.4250 to 0.4557 at @5, +7.22%; and from 0.5308 to 0.5549 at @10, +4.54%). These gains align with our design: collective perception–enhanced retrieval raises recall, and interaction chain–guided reasoning improves precision by selecting components that are experimentally suitable rather than merely topically similar.
6.3. Ablation Study
Ablation Study of Retriever.
Removing collective perception (CP) causes the largest drop. On the baseline side, adding CP over “w/o CP” yields +0.0807 Recall@20 (+21.72%), +0.0998 HitRate@5 (+21.15%), and +0.1066 HitRate@10 (+18.73%); adding self-description on top of CP still contributes (+0.0326 Recall@20, +7.77%; +0.0368 @5, +6.88%; +0.0546 @10, +8.79%). The dataset side shows the same pattern: CP accounts for the majority of the gains (+0.0508 Recall@20, +20.38%; +0.0958 @5, +29.10%; +0.1146 @10, +27.53%), while self-description brings steady, smaller improvements (+0.0292 Recall@20, +10.78%; +0.0234 @5, +5.83%; +0.0330 @10, +6.63%). This confirms that community citation contexts supply decisive third-person signals about how components are used, beyond first-person descriptions.
Ablation Study of Reranker.
Both interaction chains and supervised fine-tuning (SFT) are crucial. On the baseline side, removing interaction chains reduces HitRate@5 by 0.0688 (-13.12% relative) and HitRate@10 by 0.0583 (-9.17% relative); removing SFT yields an even larger drop (-0.1649 @5, -38.49% relative; -0.1126 @10, 19.37% relative). The dataset side mirrors this: -0.0431 @5 (-10.45% relative) and -0.0502 @10 (-9.95% relative) without chains; and -0.0916 @5 (-25.16% relative) and -0.1301 @10 (-30.63% relative) without SFT. Overall, interaction chains provide high-precision evidential priors about co-usage, and SFT teaches the model to exploit these priors to make calibrated, interpretable decisions—together delivering the observed precision gains.
Ablation Study of Recommendation Pipeline.
To assess the impact of the reranking module, we performed an ablation by removing the reranker (denoted w/o Reranker) and comparing its performance to the full model. As shown in Table 5, removing the reranker caused a notable drop in HitRate@5 and HitRate@10 for both the baseline and dataset recommendation tasks. This performance loss shows that while the retriever identifies candidate items, it cannot effectively prioritize the most relevant ones. In contrast, the reranking component, utilizing interaction chains and contextual reasoning, refines rankings and improves both ranking quality and recommendation precision. These results highlight the importance of integrating retrieval and reranking, demonstrating that the reranker significantly enhances the recommendation pipeline’s effectiveness.
@@ 6.3. Ablation Study
AgentExpt(full) 0.3001 0.4250 0.5308
Table 3. Ablation study of retriever
Method Hitrate@5 Hitrate@10
@@
Table 4. Ablation study of reranker
Method Hitrate@5 Hitrate@10

### v3

@@ 4.2 Automated Dataset Construction Pipeline
We systematically acquire targeted research papers and their corresponding metadata from prominent AI academic repositories, including OpenReview (n.d.), ArXiv (n.d.), and official conference websites. The collected metadata includes comprehensive information such as paper titles, authors, abstracts, publication years, conference affiliations, and direct links to publications. After rigorous de-duplication and cleaning, the corpus exceeding 108,825 papers was refined into a high-quality dataset suitable for subsequent analyses. Utilizing the MinerU tool (Wang et al., 2024), we perform structured parsing on each paper, accurately extracting chapter structures, paragraph boundaries, table arrangements, and citation anchors, thereby ensuring reliable downstream entity extraction.
@@
To identify baselines and datasets used in each research, we resort to citations within experiment-centric sections (e.g., Experiments, Results, Evaluation) as anchors. For baseline identification, we analyze the citation context combined with additional metadata, such as introductory papers and repository links, that are collected from resources like Paper With Code (n.d.) and OpenAlex (Priem et al., 2022). For dataset identification, we use the DataStet (Lopez, 2024) model to automatically extract dataset entities from the experimental sections, validating them through metadata from platforms like HuggingFace (n.d.) and Kaggle (n.d.). This process includes entity normalization, alias resolution, and filtering, ensuring high-quality and accurate data for subsequent validation stages.
Figure 3: The reuse rate of experiment baselines and datasets. The vertical axis indicates the reuse rate, calculated as the fraction of resources employed in year N that were introduced in preceding years (1 to N-1), reflecting the dependency on established experimental components over time.
Figure 4: Illustration of Collective Perception Augmented Retrieval
@@
To further enhance extraction accuracy and effectively mitigate false-positive candidates, we introduce a two-stage filtering strategy. Initially, we employ rule-based filtering mechanisms to preliminarily evaluate candidates by leveraging criteria such as contextual positioning, citation frequency, and naming consistency. Subsequently, we introduce an additional verification stage for entities near the decision boundary using few-shot prompting with GPT-4o. Specifically, we implement a unified context-aware evaluation protocol where the LLM integrates extracted textual evidence with external metadata to perform entity normalization and rigorous verification, ensuring high-confidence identification for both baselines and datasets.
4.3 Data Analysis
We conduct an extensive quality assessment and statistical analysis of the curated knowledge base. It comprises 116,970 distinct baseline methods and 68,316 evaluation datasets used by 108,825 research papers, establishing approximately 326,697 paper-dataset links and 543,205 paper-baseline links. Additionally, the completeness ratio of repository links achieves 13%.
A key finding from our analysis is the remarkably high reuse rate of established resources within the community. As Figure 3 shows, both baselines and datasets exhibit strong recurring adoption patterns over time. For instance, the baseline reuse rate consistently remains above 70% in recent years, with several periods even exceeding the 85% threshold. This indicates that our dataset successfully captures the core set of canonical methods and benchmarks that form the foundation for experimental comparisons in AI research.
@@ 4.3 Data Analysis
5 Method
We design two modules for dataset and baseline recommendation. (1) Collective perception-augmented retrieval: We represent each candidate by concatenating its self-description with collective perception, and finetune an embedding model on these representations to improve recall of components. (2) Reasoning-Augmented Reranking: We extract paper-baseline–paper-dataset (paper-dataset–paper-baseline) interaction chains from the literature and train a language model to generate reasoning chains over these interactions to rerank the shortlist.
5.1 Collective Perception-Augmented Retrieval
We build retrieval representations that jointly encode a candidate’s first-person self-description and third-person collective perception distilled from citation contexts, and then finetune a dense retriever to maximize recall of baselines and datasets, as shown in Figure 4.
@@ 5.1 Collective Perception-Augmented Retrieval
Our goal is to capture a target’s collective perception. A naive choice is to treat the pooled citation contexts $C(t)$ as this signal. We avoid this because a popular target may accrue thousands of mentions, far beyond the embedding model’s context window. To address these issues, We instantiate a target’s Collective Perception (CP) as a citation context summary. For each target $t$, we use GPT-4o (OpenAI, 2024) to synthesize $\mathrm{CP}(t)$ from its pooled citation contexts $C(t)$. The summary answers when and why the baseline or dataset is used, distilling recurring signals about task assumptions, evaluation protocols (metrics, splits, scales), typical configurations, and common co-usage patterns. During synthesis, we de-duplicate paraphrases, discard background or non-experimental mentions, and enforce an evidence-grounded prompt that favors majority-supported observations and avoids speculation.
@@
Figure 5: Constructing chain-derived candidates. From the interaction graph of papers, baselines, and datasets, we extract interaction chains and aggregate the terminal items to form a chain-derived dataset/baseline pool.
Table 2: Comparison on paper–baseline and paper–dataset recommendation with conservatively estimated Recall@10/20/30 and Hitrate@5/10/15.
Paper–Baseline Paper–Dataset
Method R@10 R@20 R@30 HR@5 HR@10 HR@15 R@10 R@20 R@30 HR@5 HR@10 HR@15
DataHunter (Färber and Leisinger, 2021a) 0.1809 0.3105 0.3947 0.3913 0.4862 0.6025 0.1302 0.2150 0.2444 0.3091 0.4378 0.4839
DataFinder (Viswanathan et al., 2023) 0.2918 0.4168 0.4793 0.5360 0.6216 0.6730 0.1803 0.2575 0.2961 0.3912 0.4778 0.5298
@@
where $\mathcal{T}$ describes the task (e.g., “Given a research idea, retrieve relevant baseline methods or datasets that are most suitable.”). Candidates remain in their native form without an instruction:
@@
$\mathcal{L}=\underbrace{\frac{1}{B}\sum_{i=1}^{B}\mathrm{BCE}~\!\big(S_{ii},\,y_{i}\big)}_{\text{diagonal (main)}}\;+\;\lambda\,\underbrace{\frac{1}{B(B-1)}\sum_{i\neq j}\mathrm{BCE}~\!\big(S_{ij},\,0\big)}_{\text{in-batch contrastive}}.$
@@
5.2 Reasoning-Augmented Reranking
This stage constructs reasoning chain representations from observed P-B-P-D/P-D-P-B interactions(see Figure 5) and finetunes a language model to rerank the shortlist with interpretable, chain-based justifications.
@@ 5.2 Reasoning-Augmented Reranking
which supports baseline recommendation via $p\!\to\!d\!\to\!p^{\prime}\!\to\!b$. Dataset recommendation is symmetric: $p\!\to\!b\!\to\!p^{\prime\prime}\!\to\!d$. We provide a detailed analysis validating the high coverage and selectivity of these chains in Appendix A.5.
@@
To unlock LLM reasoning and enable reasoning reranking based on interaction chains, we use GPT-4o to generate reasoning chain trajectories. For each query and each candidate baseline $b$, we enumerate chains of the form $p\!\to\!d\!\to\!p^{\prime}\!\to\!b$ and score each chain by
@@
$m_{B\!\to D}(b,d)\;=\;\bigl|\{\,p^{\prime\prime}\;:\;p^{\prime\prime}\text{ uses }b\text{ and }d\,\}\bigr|$
@@
6 Experiment
6.1 Experimental Setup
@@
To strictly evaluate performance, we compare our method against seven representative baselines spanning four distinct paradigms: (i) Lexical Matching: BM25 (Robertson et al., 2009); (ii) Supervised Classification: DataHunter (Färber and Leisinger, 2021a); (iii) Dense Retrieval: SciBERT (Beltagy et al., 2019) and the state-of-the-art retriever DataFinder (Viswanathan et al., 2023); and (iv) Structure and Graph-aware Methods: Textual-GCL (Qayyum et al., 2025), HAtten (Gu et al., 2022), and SymTax (Goyal et al., 2024). Detailed descriptions for these baselines are provided in Appendix A.1.
@@
We evaluate our framework using a rigorous chronological split (8:1:1 for train/val/test) to prevent temporal leakage and simulate real-world usage. For retrieval, we fine-tune Qwen3-Embedding-0.6B to recall the top-20 candidates. For the reasoning-augmented reranking stage, we fine-tune DeepSeek-R1-Distill-Qwen-7B to reorder these candidates. Further training configurations and hardware specifications are detailed in Appendix A.1.
6.2 Main Results
As detailed in Table 2, AgentExpt consistently outperforms all baseline methods across both recommendation tasks. Compared to the strongest prior competitor, SymTax (Goyal et al., 2024), our framework yields substantial improvements in both retrieval coverage and ranking precision. Specifically, for paper-baseline recommendation, AgentExpt improves Recall@20 by 7.23% and HitRate@10 by 7.52%. Similarly, in the paper-dataset setting, our method surpasses SymTax by 4.46% in Recall@20 and notably by 8.27% in HitRate@10. It is worth highlighting that the improvements in precision-oriented metrics (HitRate) generally exceed those in Recall, particularly for dataset recommendation. This trend validates our two-stage design: while Collective Perception ensures a high-quality candidate pool, the Reasoning-Augmented Reranker effectively leverages interaction chains to prioritize experimentally suitable components over those that are merely topically similar.
Table 3: Ablation study of retriever.
@@ 6.2 Main Results
AgentExpt (Retriever) 0.3001 0.4250 0.5308
6.3 Ablation Study
Ablation of Retriever.
As Table 3 shows, both collective perception and self-description contribute vital signals to retrieval process. Removing collective perception causes the most substantial performance degradation, with Recall@20 dropping by approximately 17.8% for baselines and 16.9% for datasets. Excluding self-descriptions also leads to a consistent decline, reducing Recall@20 by 7.2% and 9.7% respectively. These findings indicate that while first-person descriptions provide a necessary foundation, the third-person insights derived from community citation contexts are decisive for accurately capturing experimental suitability.
Ablation of Reranker.
As Table 4 shows, both the interaction chains and the reasoning augmentation contribute significantly to final performance. Removing the reasoning supervision causes a sharp decline in HitRate@5, dropping by 27.8% for baselines and 20.1% for datasets. Similarly, excluding interaction chains leads to performance degradation, with HitRate@5 falling by 11.6% and 9.5% respectively. These results confirm that high-quality evidence (interaction chains) and the capability to interpret it (reasoning alignment) are both indispensable for accurate reranking.
Ablation of Recommendation Pipeline.
To assess the impact of the reranking module, we perform an ablation by removing the reranker (denoted w/o Reranker) and comparing its performance to the full model. As Table 5 shows, removing the reranker caused a notable drop in HitRate@5 and HitRate@10 for both the baseline and dataset recommendation tasks. This performance loss shows that while the retriever identifies candidate items, it cannot effectively prioritize the most relevant ones. In contrast, the reranking component, utilizing interaction chains and contextual reasoning, refines rankings and improves both ranking quality and recommendation precision. These results highlight the importance of integrating retrieval and reranking, demonstrating that the reranker significantly enhances the recommendation pipeline’s effectiveness.
Table 4: Ablation study of reranker.
Method HR@5 HR@10
@@ 6.3 Ablation Study
Table 5: Ablation study of recommendation pipeline.
Method HR@5 HR@10